Given that the equation of tangent at the point (at^2 , 2at) on the parabola y^2 = 4ax is x - ty + at^2 = 0
A(at1^2 , 2at1) and B(at2^2 , 2at2) are points on this parabola. The equation of the chord AB is 2x - (t1 + t2)y + 2at1at2 = 0
The tangents A and B meet at the point P. Find the coordinates of P.
The line through P parallel to the axis of the parabola meets the chord AB at M. Find the coordinates of M.
1
Expert's answer
2021-02-02T05:01:17-0500

The tangent A


"x - t_1y + at_1^2 = 0"

The tangent B


"x - t_2y + at_2^2 = 0"

The tangents A and B meet at the point P


"t_1y - at_1^2 =t_2y - at_2^2"

Since "t_1\not=t_2"


"y=\dfrac{a(t_2^2-t_1^2)}{t_2-t_1}"

"y=a(t_2+t_1)"

"x=\dfrac{y^2}{4a}=\dfrac{a(t_1+t_2)^2}{4}"

"P\big(\dfrac{a(t_1+t_2)^2}{4}, a(t_2+t_1)\big)"


The equation of the axis of the parabola: "y=0."

The equation of the line through P parallel to the axis of the parabola:

"y=a(t_1+t_2)"

The equation of the chord AB is


"2x-(t_1+t_2)y+2at_1t_2=0"

The line through P parallel to the axis of the parabola meets the chord AB at M


"2x-(t_1+t_2)(a(t_1+t_2))+2at_1t_2=0"

"2x=a(t_1^2+2at_1t_2+t_2^2-2t_1t_2)"

"x=\dfrac{1}{2}(at_1^2+at_2^2))"

"M\big(\dfrac{1}{2}(at_1^2+at_2^2)), \dfrac{1}{2}(2at_1+2at_2))\big)"


Therefore M is the midpoint of AB.